OpenStudy (3psilon):

If f(2x) = x/3 , then what does f(x) = ?

Parth (parthkohli):

\[\rm {x \over 3} \div 2 \]

OpenStudy (anonymous):

f(2x) = x/3 f(2x/2) = (x/2)/3

OpenStudy (3psilon):

Why divide by 2 @ParthKohli

OpenStudy (anonymous):

Because 2x/2 = x. You get left with f(x).

Parth (parthkohli):

Erm, no, my reasoning is... that f(2x) is just the double of f(x).

OpenStudy (3psilon):

I like that reason @ParthKohli :) Thank you all

Parth (parthkohli):

I don't think that my reasoning always works, but it does in this case.

Parth (parthkohli):

Perhaps I should call it, “that reasoning.” lol

Parth (parthkohli):

It's not like I have copyrighted everything.

Parth (parthkohli):

It's also not like I'm talking to myself right now.

OpenStudy (3psilon):

lolol

OpenStudy (anonymous):

Good one,parth!!!

OpenStudy (anonymous):

Divide by 2 on each side.